OpenStudy (madmerc):

Using the completing-the-square method, rewrite f(x) = x2 + 4x − 1 in vertex form.

OpenStudy (anonymous):

nah do it the easy way

OpenStudy (madmerc):

@satellite73 do you know how to do it?

OpenStudy (anonymous):

there is a much easier way that "completing the square" not that it is not useful to know

OpenStudy (madmerc):

the choices are f(x) = (x + 2)2 + 1 f(x) = (x + 2)2 f(x) = (x + 2)2 + 4 f(x) = (x + 2)2 − 5

OpenStudy (anonymous):

yes i know how to do it the first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\) in your case \(a=1,b=4\) and \(-\frac{b}{2a}=-2\)

OpenStudy (anonymous):

so you know it will be \[f(x)=(x+2)^2+\text{something}\]

OpenStudy (anonymous):

to find the "something" replace \(x\) by \(-2\) and compute \[y=(-2)^2+4\times (-2)+1\]

OpenStudy (taylor0402):

look here im going to give you an example Example: y = x^2 + 4x - 6 You have to complete the square to convert it into vertex form. y = x² + 4x - 6 Group. y = (x² + 4x) - 6 Factor y = (x² + 4x) - 6 Add placeholders. y = (x² + 4x + ___) - 6 - 1(___) Notice that the second blank is multiplied by -1 to account for what you had to add to complete the square. Take the coefficient of the x term: 4 Divide it by 2: 4 / 2 = 2 Square it: (2)² = 4 Add 4 to both blanks. y = (x² + 4x + 4) - 6 - 1(4) x² + 4x + 4 is the expanded form of a perfect square binomial. Remember that (a + b)² = a² + 2ab + b². Apply this to what you have. y = (x² + 4x + 4) - 6 - 1(4) y = (x + 2)² - 6 - 1(4) Simplify the rest. y = (x + 2)² - 6 - 4 y = (x + 2)² - 10 Remember that the vertex form is: y = a(x - h)² + k CHECK: y = (x + 2)² - 10 y = [(x)² + 2(x)(2) + (2)²] - 10 y = (x² + 4x + 4) - 10 y = (x²) + 1(4x) + 1(4) - 10 y = x² + 4x + 4 - 10 y = x² + 4x - 6 TRUE ANSWER: y = (x + 2)² - 10 is the vertex form. Given: y = (x + 2)² - 10 Means: h = -2 Means: k = -10 Means: a = 1 ANSWER: The vertex is at (-2, -10). Since the equation is a function of x and a is positive, the parabola opens upwards. Parabolas that open upwards or downwards have an axis of symmetry that is the same as the h coordinate of the vertex. ANSWER: The axis of symmetry is at x = -2. Remember that y-intercepts have x = 0. Find the value of y when x = 0 by substitution x with 0 in the original equation. y = x² + 4x - 6 y = 0² + 4(0) - 6 y = 0 + 0 - 6 y = -6 ANSWER: The y-intercept is at (0, -6).

OpenStudy (anonymous):

that is the number out at the end

OpenStudy (taylor0402):

is that helpful 0_o

OpenStudy (madmerc):

wait so the answer to my question would be f(x) = (x + 2)^2 + 1 ?

OpenStudy (madmerc):

@Taylor0402 @satellite73

OpenStudy (anonymous):

hold on let me check

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[(-2)^2+4\times (-2)+1\] is not \(1\)

OpenStudy (madmerc):

I got -7 for the answer to that @satellite73

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

\[(-2)^2+4\times (-2)+1\neq -7\]

OpenStudy (madmerc):

is it -3? i dont know how im messing this up @satellite73

OpenStudy (anonymous):

\[4-8+1=-4+1=-3\]

OpenStudy (anonymous):

oooh hold on, typeo!!

OpenStudy (anonymous):

it is \[x^2+4x-1\] so \[4-8-1=-5\]

OpenStudy (anonymous):

i put plus one, it should have been minus one, my fault

OpenStudy (madmerc):

oooh okay! so it's f(x) = (x + 2)^2 − 5

OpenStudy (madmerc):

can you help me with another one? @satellite73

OpenStudy (anonymous):

sure

OpenStudy (madmerc):

Solve x2 − 12x + 5 = 0 using the completing-the-square method. @satellite73

OpenStudy (anonymous):

ok this time we really will complete the square, because it is easiest done that way

OpenStudy (anonymous):

first subtract 5 from both sides to get \[x^2-12x=-5\] now, what is half of 12?

OpenStudy (madmerc):

6 @satellite73

OpenStudy (anonymous):

right and \(6^2=?\)

OpenStudy (madmerc):

36

OpenStudy (anonymous):

exactly so we go from \[x^2-12x=-5\] to \[(x-6)^2=-5+36\] or \[(x-6)^2=31\] that is called "completing the square"

OpenStudy (anonymous):

now we are ready to solve it \[(x-6)^2=31\\ x-6=\pm\sqrt{31}\\ x=6\pm\sqrt{31}\]

OpenStudy (madmerc):

thank you so much!!!! can you help me with one more? @satellite73

OpenStudy (anonymous):

ok but this one will cost you $7

OpenStudy (anonymous):

just kidding

OpenStudy (madmerc):

thank you!!!! okay its Rewrite f(x) = –2(x − 3)2 + 2 from vertex form to standard form. i dont know the difference b/w vertex and standard

OpenStudy (anonymous):

it just means multiply out and combine like terms to make it look like \[f(x)=ax^2+bx+c\]

OpenStudy (anonymous):

i.e. write out \[-2(x-3)(x-3)+2\] multiply, distribute the \(-2\) etc

OpenStudy (madmerc):

@satellite73 okay so, f(x) = 4x^2 - 24x + 38 ?

OpenStudy (anonymous):

that cannot be right

OpenStudy (anonymous):

don't square the \(-2\)

OpenStudy (anonymous):

square the \(x-3\) i.e. compute \((x-3)(x-3)\) then multipiy the result by \(-2\)

OpenStudy (madmerc):

f(x) = -2x + 12x - something? @satellite73

OpenStudy (madmerc):

f(x) = –2x2 + 12x − 16 !!! right?

OpenStudy (madmerc):

sorry -2x^2 + 12x - 16